# Intermediate mass-ratio inspirals in a dense dark-matter environment: Effects of the initial dark-matter distribution

### v1

@@ IV.2 Gravitational-wave dephasing
where $m_{2,0}=10\,\mathrm{M}_{\odot}$ is the initial mass at $3r_{\mathrm{2,4y}}^{\mathrm{A}}$ and $\delta m_{\mathrm{2,4y}}^{\mathrm{A}}$ is defined as the difference $m_{\mathrm{2,4y}}^{\mathrm{A}}-m_{2,0}$. Specifically, we list $\delta m_{\mathrm{2,4y}}^{1}$ and $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{I}}}$ in Table 1. The results for $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ are similar to those in [48]. However, because we use a somewhat lower density of $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$ here and include the evolution of the mass $m_{2}$ in the equations of motion for $r_{2}$, the two cases are not directly comparable. The trend in $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ as a function of $m_{1}$, and the fact that the largest absolute and fractional dephasing against vacuum with $j_{\mathrm{min}}=0$ happens with a primary mass of $m_{1}=3\times 10^{4}~\mathrm{M}_{\odot}$, is the same as in [48]. Therefore, we refer to [48] for the discussion of this trend. Here we focus on any differences in the dephasings $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ and $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{I})}}$ in the third and sixth columns, respectively, of Table 1.
@@
The values of $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ and $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{I})}}$ were computed from different initial radii (starting from $r_{\mathrm{2,4y}}^{\mathrm{A}}$ in each case, respectively), which should be considered when comparing these values for the same primary mass $m_{1}$. Nevertheless, we can identify general trends in the effect of introducing the angular-momentum cutoff on the dephasing against vacuum for a given $m_{1}$. The largest absolute difference in dephasing between simulations with $j_{\mathrm{min}}=\sqrt{8}$ and $j_{\mathrm{min}}=0$ occurs for two values: $m_{1}=3\times 10^{4}\,\mathrm{M}_{\odot}$ and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. Given the smaller number of cycles for $m_{1}=10^{5}\,\mathrm{M}_{\odot}$, the largest fractional difference occurs for $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. As the mass ratio becomes more extreme, the secondary orbits at a smaller number of gravitational radii during the last four years of the inspiral than it does for less extreme mass ratios. Because the densities with $j_{\mathrm{min}}=\sqrt{8}$ differ most from those with $j_{\mathrm{min}}=0$ near the ISCO, it is reasonable that a nonzero $j_{\mathrm{min}}$ has the largest effect for these more extreme mass ratios. That this difference is the same at both $m_{1}=3\times 10^{4}\,\mathrm{M}_{\odot}$ and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ is likely coincidental. However, the trend that the total number of orbital cycles between $r_{\mathrm{2,4y}}^{\mathrm{A}}$ and ISCO decreases as the mass ratio becomes more extreme means the number of cycles of dephasing also decreases.
@@
In [48], it was found that accretion between $r_{\mathrm{2,4y}}$ and $r_{\mathrm{ISCO}}$ increased as the mass ratio became less extreme, whereas here we find that both $\delta m_{\mathrm{2,4y}}^{1}$ and $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{I}}}$ (the accretion that takes place between $3r_{\mathrm{2,4y}}$ and $r_{\mathrm{2,4y}}^{\mathrm{A}}$) is largest for $m_{1}=3\times 10^{3}\mathrm{M}_{\odot}$ and decreases for both less and more extreme mass ratios. Although there is no reason that in different spatial regions the accretion should behave similarly as a function of mass ratio, one might have anticipated a common trend. Unlike the the accretion between $r_{\mathrm{2,4y}}$ and $r_{\mathrm{ISCO}}$ in [48], the values of $\delta m_{\mathrm{2,4y}}^{\mathrm{A}}$ are affected by the initial conditions at $3r_{\mathrm{2,4y}}$ that are inconsistent with an adiabatic inspiral from larger radii. Thus, we do not dwell on the details of the $m_{1}$ dependence of the results for $\delta m_{\mathrm{2,4y}}^{\mathrm{A}}$ given the unrealistic initial conditions that necessarily arise in their computation. Nevertheless, we speculate that the $\delta m_{\mathrm{2,4y}}^{\mathrm{A}}$ are smaller for $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ than for $m_{1}=3\times 10^{3}\,\mathrm{M}_{\odot}$, because DF feedback and SA feedback are more efficient for the $10^{3}\,\mathrm{M}_{\odot}$ case, and because the $10^{3}\,\mathrm{M}_{\odot}$ case starts at a larger number of gravitational radii from $m_{1}$, where the corresponding densities are lower.
@@
We conclude this subsection with a remark about the choice of the starting separation $r_{2}=3r_{\mathrm{2,4y}}$, which we selected to obtain reasonable initial conditions for the density at $r_{2}=r_{\mathrm{2,4y}}^{\mathrm{A}}$, the largest radius where we compute the dephasing against vacuum binaries. One might wonder whether introducing the cutoff in angular momentum has any effect on this starting separation, which was determined in [40, 48] to be an appropriate choice for $j_{\mathrm{min}}=0$. [^8] Specifically, in [40, 48] it was noted that, when $r_{2}=3r_{\mathrm{2,4y}}$, the secondary’s influence on the DM distribution was small at radii smaller than $2r_{\mathrm{2,4y}}$, so that the density at $r=r_{\mathrm{2,4y}}$ was largely unaffected. For this reason, initializing the secondary at $r_{2}=3r_{\mathrm{2,4y}}$ was taken to be consistent with an adiabatic inspiral from further out. We performed a somewhat different test here. Specifically, we simulated inspirals starting from $r_{2}=2r_{\mathrm{2,4y}}$ and $r_{2}=4r_{\mathrm{2,4y}}$, and we compared the dephasing from these simulations against those given in Table 1. The magnitude of the change in dephasing depends on the mass ratio. Nevertheless, the difference in dephasing, when comparing the simulations starting at $r_{2}=4r_{\mathrm{2,4y}}$ against those at $r_{2}=3r_{\mathrm{2,4y}}$, was $\mathcal{O}(10)$ cycles or fewer, for the cases that we considered (again below the accuracy at which we present our results). For a starting separation of $r_{2}=2r_{\mathrm{2,4y}}$, there were differences in the dephasing of as much as a few hundred cycles, which is larger than the accuracy of our results. Thus, the starting separation of $3r_{\mathrm{2,4y}}$ is also appropriate for the cutoff of $j_{\mathrm{min}}=\sqrt{8}$.
@@ IV.3 Dark-matter density
Figure 3: Dark-matter density at three times during an inspiral. In all the times depicted, the density has a cutoff in the angular momentum ($j_{\mathrm{min}}=\sqrt{8}$) so that it vanishes at $r=2R_{\mathrm{s}}$. We chose $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$ and $\gamma_{\mathrm{sp}}=7/3$ for the dark-matter density and $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ for the primary mass. The dotted black line is the initial density profile (when $r_{2}=3r_{\mathrm{2,4y}}$, but before the evolution begins), the dash-dotted orange curve is the density when $r_{2}=r_{\mathrm{2,4y}}$, and the solid blue is the density when the secondary reaches the ISCO. Further discussion of the figure is given in Sec. IV.3.
@@
Figure 3 shows the dark-matter density at three stages during the inspiral of a $m_{1}=10^{3}\,\mathrm{M}_{\odot}$, $m_{2,0}=10\,\mathrm{M}_{\odot}$ binary. The initial density is the dotted black curve. This corresponds to a binary separation of $r_{2}$ = $3r_{\mathrm{2,4y}}\approx 4.9\times 10^{-8}\,\mathrm{pc}$, but the binary has not yet begun its inspiral. The density for a binary separation of $r_{2}$ = $r_{\mathrm{2,4y}}$ is the dash-dotted orange line, and the solid blue curve is the density when the secondary reaches the ISCO (the end of our simulation). The location of the secondary at $r_{2}$ = $r_{\mathrm{2,4y}}$ is marked with an orange star.
@@
The combination of DF and SA feedback has a significant effect on the dark-matter density. Throughout the inspiral, feedback from dynamical friction depletes the density of particles near the secondary by scattering them to larger radii. Secondary accretion leads to a decrease in the density by removing DM particles from the distribution function. The decrease in the density between the initial time and final time in the simulation (the ISCO case) is around an order of magnitude at radii smaller than $3r_{\mathrm{2,4y}}$, and somewhat larger at the larger end of this interval. At radii larger than $r_{\mathrm{2,4y}}$, the densities look qualitatively similar to those with $j_{\mathrm{min}}=0$ (see the right panel of Fig. 6 in [48]). However, there are more significant differences between the $j_{\mathrm{min}}=0$ in [48] and the $j_{\mathrm{min}}=\sqrt{8}$ for radii less than $r_{\mathrm{2,4y}}$.
@@
In Fig. 3, the density after the inspiral (the ISCO case) decreases by almost an order of magnitude for $r\leq 10^{-9}$ pc, whereas it was closer to a factor of a few for the $j_{\mathrm{min}}=0$ given in [48]. As in the $j_{\mathrm{min}}=0$ case, DF feedback has a larger impact on the density than SA feedback for this region as a whole. Specifically, the mass enclosed within $r_{\mathrm{2,4y}}$ is initially roughly $5\times 10^{-3}\,\mathrm{M}_{\odot}$, and it changes by roughly $4\times 10^{-3}\,\mathrm{M}_{\odot}$ between the start of the simulation and the end. We can compute how much of the mass is accreted by the secondary (about $30\%$) and how much is ejected from the spike (about $10\%$). The remaining amount ($60\%$ of the mass decrease) arises from the scattering particles from smaller radii to larger radii, which is characteristic of dynamical friction.
@@ V Simulations of successive mergers
We briefly discuss why this is a probable merger scenario. First, IMBHs may form in a variety of environments, including within globular clusters or at the center of dwarf galaxies (see, e.g., [33]). These environments can host a population of stellar-mass black holes in proximity to the IMBH. For clumpy star-forming clusters with a central IMBH surrounded by stellar-mass objects (stars and compact objects), [34] predicts an IMRI rate of up to $10^{-7}$ yr-1, while [35] predicts up to $10^{-8}$ yr-1 for IMBHs in local-universe globular clusters. Each IMRI merger will modify the DM distribution around the IMBH through DF and SA feedback. [^9] Each merger will also slightly spin up the primary black hole by an amount of order $qGm_{1}^{2}/c$; thus, it would take of order $q^{-1}$ mergers to appreciably spin up the primary if it were initially nonspinning. We continue to ignore the spin of the primary, as a result. The predicted time between mergers ($10^{7}$$10^{8}$ years) is shorter than the characteristic relaxation time of the DM distribution (which had been estimated in [40] to be many orders of magnitude larger). This makes it a reasonable approximation to assume that the DM distribution does not evolve from the post-merger configuration following the prior IMRI merger.

### v2

@@ IV.2.2 Discussion of dephasing results
We present in Table 1 the cycles and dephasing against vacuum for the five simulations described in Sec. IV.1. We also list $\delta m_{\mathrm{2,4y}}^{(1)}$ and $\delta m_{\mathrm{2,4y}}^{\mathrm{(2_{I})}}$ in Table 1. The results for $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ are similar to those in [53]. However, because we use a somewhat lower density of $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$ here and include the evolution of the mass $m_{2}$ in the equations of motion for $r_{2}$, the two cases are not directly comparable. The trend in $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ as a function of $m_{1}$, and the fact that the largest absolute and fractional dephasing against vacuum with $j_{\mathrm{min}}=0$ happens with a primary mass of $m_{1}=3\times 10^{4}~\mathrm{M}_{\odot}$, is the same as in [53]. Therefore, we refer to [53] for the discussion of this trend. Here we focus on any differences in the dephasings $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ and $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{I})}}$ in the third and sixth columns, respectively, of Table 1.
@@
The values of $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-1)}}$ and $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{I})}}$ were computed from different initial radii (starting from $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$ in each case, respectively), which should be considered when comparing these values for the same primary mass $m_{1}$. Nevertheless, we can identify general trends in the effect of introducing the angular-momentum cutoff on the dephasing against vacuum for a given $m_{1}$. The largest absolute difference in dephasing between simulations with $j_{\mathrm{min}}=\sqrt{8}$ and $j_{\mathrm{min}}=0$ occurs for two values: $m_{1}=3\times 10^{4}\,\mathrm{M}_{\odot}$ and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. Given the smaller number of cycles for $m_{1}=10^{5}\,\mathrm{M}_{\odot}$, the largest fractional difference occurs for $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. As the mass ratio becomes more extreme, the secondary orbits at a smaller number of gravitational radii during the last four years of the inspiral than it does for less extreme mass ratios. Because the densities with $j_{\mathrm{min}}=\sqrt{8}$ differ most from those with $j_{\mathrm{min}}=0$ near the ISCO, it is reasonable that a nonzero $j_{\mathrm{min}}$ has the largest effect for these more extreme mass ratios. That this difference is the same at both $m_{1}=3\times 10^{4}\,\mathrm{M}_{\odot}$ and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ is likely coincidental. However, the trend that the total number of orbital cycles between $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$ and ISCO decreases as the mass ratio becomes more extreme means the number of cycles of dephasing also decreases.
@@
In [53], it was found that accretion between $r_{\mathrm{2,4y}}$ and $r_{\mathrm{ISCO}}$ increased as the mass ratio became less extreme, whereas here we find that both $\delta m_{\mathrm{2,4y}}^{(1)}$ and $\delta m_{\mathrm{2,4y}}^{\mathrm{(2_{I})}}$ (the accretion that takes place between $3r_{\mathrm{2,4y}}$ and $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$) is largest for $m_{1}=3\times 10^{3}\mathrm{M}_{\odot}$ and decreases for both less and more extreme mass ratios. Although there is no reason that in different spatial regions the accretion should behave similarly as a function of mass ratio, one might have anticipated a common trend. Unlike the the accretion between $r_{\mathrm{2,4y}}$ and $r_{\mathrm{ISCO}}$ in [53], the values of $\delta m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ are affected by the initial conditions at $3r_{\mathrm{2,4y}}$ that are inconsistent with an adiabatic inspiral from larger radii. Thus, we do not dwell on the details of the $m_{1}$ dependence of the results for $\delta m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ given the unrealistic initial conditions that necessarily arise in their computation. Nevertheless, we speculate that the $\delta m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ are smaller for $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ than for $m_{1}=3\times 10^{3}\,\mathrm{M}_{\odot}$, because DF feedback and SA feedback are more efficient for the $10^{3}\,\mathrm{M}_{\odot}$ case, and because the $10^{3}\,\mathrm{M}_{\odot}$ case starts at a larger number of gravitational radii from $m_{1}$, where the corresponding densities are lower.
@@
We conclude this subsection with a remark about the choice of the starting separation $r_{2}=3r_{\mathrm{2,4y}}$, which we selected to obtain reasonable initial conditions for the density at $r_{2}=r_{\mathrm{2,4y}}^{\mathrm{(A)}}$, the largest radius where we compute the dephasing against vacuum binaries. One might wonder whether introducing the cutoff in angular momentum has any effect on this starting separation, which was determined in [46, 53] to be an appropriate choice for $j_{\mathrm{min}}=0$. [^11] Specifically, in [46, 53] it was noted that, when $r_{2}=3r_{\mathrm{2,4y}}$, the secondary’s influence on the DM distribution was small at radii smaller than $2r_{\mathrm{2,4y}}$, so that the density at $r=r_{\mathrm{2,4y}}$ was largely unaffected. For this reason, initializing the secondary at $r_{2}=3r_{\mathrm{2,4y}}$ was taken to be consistent with an adiabatic inspiral from further out. We performed a somewhat different test here. Specifically, we simulated inspirals starting from $r_{2}=2r_{\mathrm{2,4y}}$ and $r_{2}=4r_{\mathrm{2,4y}}$, and we compared the dephasing from these simulations against those given in Table 1. The magnitude of the change in dephasing depends on the mass ratio. Nevertheless, the difference in dephasing, when comparing the simulations starting at $r_{2}=4r_{\mathrm{2,4y}}$ against those at $r_{2}=3r_{\mathrm{2,4y}}$, was $\mathcal{O}(10)$ cycles or fewer, for the cases that we considered (again below the accuracy at which we present our results). For a starting separation of $r_{2}=2r_{\mathrm{2,4y}}$, there were differences in the dephasing of as much as a few hundred cycles, which is larger than the accuracy of our results. Thus, the starting separation of $3r_{\mathrm{2,4y}}$ is also appropriate for the cutoff of $j_{\mathrm{min}}=\sqrt{8}$.
@@ IV.3 Dark-matter density
Figure 3: Dark-matter density at three times during an inspiral. In all the times depicted in both panels, the density has a cutoff in the angular momentum ($j_{\mathrm{min}}=\sqrt{8}$) so that it vanishes at $r=2R_{\mathrm{s}}$. We chose $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$ and $\gamma_{\mathrm{sp}}=7/3$ for the dark-matter density in both panels and $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ (left), or $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ (right), for the primary mass. The dotted black line is the initial density profile (when $r_{2}=3r_{\mathrm{2,4y}}$, but before the evolution begins), the dash-dotted orange curve is the density when $r_{2}=r_{\mathrm{2,4y}}$, and the solid blue is the density when the secondary reaches the ISCO. Further discussion of the figure is given in Sec. IV.3.
@@
Figure 3 shows the dark-matter density at three stages during the inspiral of a binary with $m_{2,0}=10\,\mathrm{M}_{\odot}$ and $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ (left) or $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ (right). The initial density in each panel is the dotted black curve. In the left panel, the initial binary separation is $3r_{\mathrm{2,4y}}\approx 5\times 10^{-8}\,\mathrm{pc}$, whereas in the right panel it is $3r_{\mathrm{2,4y}}\approx 5\times 10^{-7}$ pc. The density for a binary separation of $r_{2}$ = $r_{\mathrm{2,4y}}$ is the dash-dotted orange line, and the solid blue curve is the density when the secondary reaches the ISCO (the end of our simulation) in both panels. The location of the secondary at $r_{2}=r_{\mathrm{2,4y}}$ is marked with an orange star.
@@
The combination of DF and SA feedback has a significant effect on the DM density, and the effect is stronger for the $m_{1}=10^{3}\mathrm{M}_{\odot}$ case. Throughout the inspiral, feedback from dynamical friction depletes the density of particles near the secondary by scattering them to larger radii. Secondary accretion leads to a decrease in the density by removing DM particles from the distribution function. The decrease in the density between the initial time and final time in the simulation (the ISCO case) is around an order of magnitude at radii smaller than $3r_{\mathrm{2,4y}}$, and somewhat larger at the larger end of this interval. At radii larger than $r_{\mathrm{2,4y}}$, the densities look qualitatively similar to those with $j_{\mathrm{min}}=0$ (see the right panel of Fig. 6 in [53]). However, there are more significant differences between the $j_{\mathrm{min}}=0$ in [53] and the $j_{\mathrm{min}}=\sqrt{8}$ for radii less than $r_{\mathrm{2,4y}}$.
@@
In the left panel of Fig. 3 ($m_{1}=10^{3}\,\mathrm{M}_{\odot}$), the density after the inspiral (the ISCO case) decreases by almost an order of magnitude for $r\lesssim 10^{-9}$ pc, whereas it was closer to a factor of a few for the $j_{\mathrm{min}}=0$ given in [53]. As in the $j_{\mathrm{min}}=0$ case, DF feedback has a larger impact on the density than SA feedback for this region as a whole. Specifically, the mass enclosed within $r_{\mathrm{2,4y}}$ is initially roughly $5\times 10^{-3}\,\mathrm{M}_{\odot}$, and it changes by roughly $4\times 10^{-3}\,\mathrm{M}_{\odot}$ between the start of the simulation and the end. We can compute how much of the mass is accreted by the secondary (about $30\%$) and how much is ejected from the spike (about $10\%$). The remaining amount ($60\%$ of the mass decrease) arises from the scattering particles from smaller radii to larger radii, which is characteristic of dynamical friction.
@@
In the right panel of Fig. 3 ($m_{1}=10^{5}\,\mathrm{M}_{\odot}$), we observe that the density is not changed as significantly as it is in the $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ case. There still is a decrease in the density because of dynamical friction and accretion onto the secondary. Both dynamical friction and accretion affects are proportional to the mass ratio, so it is not surprising that they become less efficient for more-extreme mass ratios. The presence of the increased density for $r\gtrsim 4\times 10^{-7}$ pc, when the secondary reaches $r_{\mathrm{2,4y}}$ and $r_{\mathrm{ISCO}}$, occurs because dynamical friction scatters particles to larger radii, but secondary accretion is not sufficiently strong to accrete these particles (as it was in the $m_{1}=10^{3}\mathrm{M}_{\odot}$ case).
@@ V Simulations of successive mergers
We briefly discuss why this is a probable merger scenario. First, IMBHs may form in a variety of environments, including within globular clusters or at the center of dwarf galaxies (see, e.g., [58]). These environments can host a population of stellar-mass black holes in proximity to the IMBH. For clumpy star-forming clusters with a central IMBH surrounded by stellar-mass objects (stars and compact objects), [55] predicts an IMRI rate of up to $10^{-7}$ yr-1, while [38] predicts up to $10^{-8}$ yr-1 for IMBHs in local-universe globular clusters. These rates imply that in a $10^{8}$ yr time period, there is an order one probability of the IMBH capturing a second compact object and undergoing a second IMRI merger. Each IMRI merger will modify the DM distribution around the IMBH through DF and SA feedback. [^12] Each merger will also slightly spin up the primary black hole by an amount of order $qGm_{1}^{2}/c$; thus, it would take of order $q^{-1}$ mergers to appreciably spin up the primary if it were initially nonspinning. We continue to ignore the spin of the primary, as a result. The merger will also give rise to a gravitational-wave kick, which scales as $q^{2}$ [37]; thus, the linear momentum imparted to the merger remnant is also negligible in our leading-order treatment in $q$. The predicted time between mergers ($10^{7}$$10^{8}$ years) is shorter than the characteristic relaxation (or “refilling”) timescale of the DM distribution, which had been estimated in [46] to be of order $10^{70}\,$yr for a 100100\,Gev DM particle (the timescale goes inversely with the DM particle mass). This makes it a reasonable approximation to assume that the DM distribution does not evolve from the post-merger configuration following the prior IMRI merger.
@@
The data files for the DM density profiles in this section (along with those described in Sec. IV) can be downloaded from [64].